OpenStudy (anonymous):

Solve the equation by factoring. 4x2 - 17x = -4

OpenStudy (anonymous):

you gotts find x first

OpenStudy (anonymous):

yeah that's y u have to factor

OpenStudy (anonymous):

@mathslover hey can you help out

mathslover (mathslover):

\(4x^2 - 17x = -4\) This is your question. Now, add 4 both sides \(4x^2 - 17x + 4 = -4 + 4\)

mathslover (mathslover):

Can you tell me what is "-4+4" or "4-4" ?

mathslover (mathslover):

@Zen_Controversy , you there?